OpenStudy (anonymous):

Given the following equation, indicate the standard equation of the conic section it represents. 16x 2 - 300 - 25y 2 = 100 Find the asymptotes. help! I don't know how to put the equation into the form of.. x^2/a^2 - y^2/b^2=1

ganeshie8 (ganeshie8):

start with adding 300 both sides

OpenStudy (anonymous):

Oh & then you divide by 400 and you get.. 16x^2/20^2 - 25y^2/20^2 = 1. Thank you :) That is silly of me to not know.. But thanks!

ganeshie8 (ganeshie8):

yep ! good work :)

OpenStudy (anonymous):

It says.. the asymptotes +/- __blank__/__blank__x Would by answer be +/-1/1x ??

ganeshie8 (ganeshie8):

im sorry i forgot how to get asymptotes of hyperbola :(

OpenStudy (anonymous):

Your fine! But thanks for your help!

ganeshie8 (ganeshie8):

gotcha. \(\pm \frac{b}{a}x \) => \(\pm\frac{4}{5}x\)

OpenStudy (anonymous):

I see you figured it out :P & Thanks youu!!!

ganeshie8 (ganeshie8):

np... :)